OpenStudy (anonymous):

2. An object is launched straight up. The object attains a height of 107 m after 2.50 s. a. What was the initial velocity? b. When (i.e. how long after launch) will the object have a speed of 20 m/s on the way back down?

OpenStudy (anonymous):

s=ut+1/2at^2 107=u*2.5- 1/2 * 10* 2.5^2 u= 55.3m/s v= u+at when reaching maximum height v=0 0=55.3-10*t1 t1=5.53s now on the way back v=20m/s 20=10*t2 t2=2 hence t1+t2=7.53 so after 7.73 seconds from the launch the body has a velocity of 20m/s on the way down